OpenStudy (nuccioreggie):

What is the value of r in the equation 3r − 6 = 18? 6 8 18 24

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (nuccioreggie):

@AlexandervonHumboldt2

OpenStudy (mathstudent55):

Try each value in the equation and see which one works.

OpenStudy (smilie2015):

3r = 18 +16 add 16 on both sides then when u get value divide it with 3 on both sides

OpenStudy (mathstudent55):

3r - 6 = 18 means 3 times a number minus 6 equals 6. Try each number in the choices. 3 times the number. Then subtract 6. The one that equals 18 is the correct choice.

OpenStudy (smilie2015):

sorry not 16 its 6

OpenStudy (mathstudent55):

@smilie2015 Look carefully at the problem. You meant 6, not 16.

OpenStudy (mathstudent55):

You got it.

OpenStudy (smilie2015):

ohhh thanks :)

OpenStudy (michele_laino):

If I add \(6\) to both sides, I get: \[\huge 3r - 6 + 6 = 18 + 6\] please simplify

OpenStudy (nuccioreggie):

so the answer is A